\begin{thebibliography}{99}








\bibitem{mss} A. V. Mikhailov, A. B. Shabat, and V. V. Sokolov, The symmetry approach to classification of integrable equations. In What is integrability?, Springer Ser. Nonlinear Dynamics, pages 115–184. Springer, Berlin, 1991.



\bibitem{OLV} P.J. Olver, Applications of Lie Groups to Differential Equations, Second Ed., Springer-Verlag, New York, 1993. 


\bibitem{w} J. P. Wang, “A list of 1 + 1 dimensional integrable equations and their properties,” J. Nonlinear Math. Phys. 9, 213, 2002.






\bibitem{s} A. Sergyeyev, Zero curvature representation for a new fifth-order integrable system. Fundam. Prikl. Mat, 12(7):227–229, 2006.



 \bibitem{mnw2}A. V. Mikhailov, V. S. Novikov, and Jing Ping Wang, On classification of integrable non-evolutionary equations. Stud. Appl. Math., 118:419–457, 2007.


\bibitem{mnw1} A. V. Mikhailov, V. S. Novikov, and J. P.Wang, “Symbolic representation and classification of integrable systems, inAlgebraic Theory of Differential Equations edited by Mikhailov, A. V. and MacCallum, M. A. H. (Cambridge University
Press, 2009), pp. 156–216 ; e-print arXiv:0712.1972.


\bibitem{v} P. Vojcak, On complete integrability of the Mikhailov-Novikov-Wang system. J. Math. Phys. 52, 043513, 2011.



\bibitem{dkt} A. De Sole, V. G. Kac and R. Turhan , A new approach to the Lenard-Magri scheme of integrability. arxiv: 1303.3438v1.




\end{thebibliography}

\documentclass[fleqn,reqno,a4paper,12pt]{amsart}

\usepackage[utf8]{inputenc}
\usepackage{amsmath}
\usepackage{amsfonts}
\usepackage{amssymb}
\usepackage[left=2cm,top=3cm,right=2cm,bottom=3cm]{geometry}

\begin{document}

\title[ A   Fifth-Order Bi-Hamiltonian System ]{A   Fifth-Order Bi-Hamiltonian  System  }

\author[Daryoush. Talati]{Daryoush Talati}
 \email{daryoush.talati@eng.ankara.edu.tr,~~daryoush.talati@gmail.com}

\address{ No. 15, 22nd Alley, Jomhoori Eslami Boulevard, Salmas, Iran}

\newtheorem{theo}{Theorem}[section]
\newtheorem{prop}{Proposition}

\begin{abstract}

In this work, we introduced a new two component fifth-order bi-Hamiltonian system admitting the scalar Kupershmidt equation as a reduction.

\end{abstract}

\maketitle

\section*{Introduction}

To prove the integrability of an equation  suspected to be bi-Hamiltonian, one need to find an appropriate compatible pair of Hamilton operator $J$ and $K$ such that the Magri scheme

\begin{equation*}
u_{t_i} =F_i[u]=KG_i[u]=JG_{i+1}[u]  ,~ i=-1,0,1,2,3,...
\end{equation*}

constructed by  the operators contains the equation in hand. Here $F_i[u]$ are characteristics of symmetries and $G_i$ are the conserved gradients.

The scalar equations of order up to 5 are extensively classified with respect to existence of sufficiently many higher conserved densities for the existence of a formal symmetry. Unlik lower order cases, classification of  fifth order two-component evolution equations is an obstinate problem. Some completely integrable equations of this type were found by  Mikhailov, Novikov and Wang \cite{mnw1,mnw2} in the study of symbolic representation theory and  nonevolutionary equations


\begin{equation}
\left(
\begin{array}{l}
u\\ \\
v\\ \\
\end{array}\right)_{t}=
\left(
\begin{array}{cc} 
-\frac{5}{3}u_{5x} -10vv_{3x}+10uu_{3x}+ 25u_{x}u_{xx}  - 15v_{x}v_{xx}  -12u^2u_{x}\\+6v^2u_{x}+12uvv_{x}-6v^2v_{x}\\
15v_{5x} -10vu_{3x}-30uv_{3x}-35v_{x}u_{xx} +30v_{x}v_{xx} -45u_{x}v_{xx}\\ +6v^2u_{x}-6v^2v_{x} + 12uvu_{x} + 12u^2v_{x}  

\end{array}
\right) ,\label{sys1}
\end{equation}

\begin{equation}
\left(
\begin{array}{l}
u\\ 
v\\ 
\end{array}\right)_{t}=
\left(
\begin{array}{cc} 
u_{5x} + 10uu_{3x} + 25u_{x}u_{xx}+20u^2u_{x}+v^2v_{x}\\
u_{3x}v + u_{xx}v_{x}+8uvu_{x}+4u^2v_{x}
\end{array}
\right) ,\label{sys2}
\end{equation}


\begin{equation}
\left(
\begin{array}{l}
u\\ 
v\\ 
\end{array}\right)_{t}=
\left(
\begin{array}{cc} 
-\frac{1}{8}u_{5x}-2 uu_{3x}-2u_{x}u_{xx}-\frac{32}{5}u^2u_{x}+v_{x}\\
\frac{9}{8}v_{5x} + 6uv_{3x} + 6u_{x}v_{xx} + 4u_{xx}v_{x}+ \frac{32}{5}u^2v_{x}
\end{array}
\right) .\label{sys3}
\end{equation}


System (\ref{sys1}) and (\ref{sys2}) admit a reduction $v=0$ to the Kaup-Kupershmidt equation $u_t=u_{5x}+10uu_{3x}+25u_{x}u_{xx}+20u^2u_{x}$ and By setting $v=0,$ system (\ref{sys3}) reduces to the Sawada-Kotera equation   $u_t=u_{5x}+5uu_{3x}+ 5u_{x}u_{xx}+5u^2u_{x}$ (see \cite {mss,w} and references therein).

Bi-Hamiltonian structures   for  (\ref{sys2}) and  (\ref{sys3})  can be found in \cite{mnw2}. Bi-Hamiltonian structure for  system (\ref{sys1}) and Zero curviture representation for (\ref{sys2})  are discussed in \cite{v,s}.
 Very recently system (\ref{sys2}) considered by De Sole, Kac and Turhan \cite{dkt} in the study of the Lenard-Magri scheme of integrability who developed a new method based on the notion of strongly skew-adjoint differential operators, using the Lie superalgebra of variational polyvector fields. However, no two-component  completely integrable  system with  reduction $v=0$ to the kupershmidt equation is known so far. In this work, we introduce a new  system of this type whose bi-Hamiltonian structure we constructed too.

\section*{The new System}

The new fifth order bi-Hamiltonian two-component system we introduce here is

\begin{equation}
\left(
\begin{array}{l}
u\\\\\\\\\\\\\\
v\\\\\\\\\\\\
\end{array}\right)_{t}=
\left(
\begin{array}{cc} 
u_{5x}-30vv_{4x}+5u_{x}u_{3x}-5u^{2}u_{3x}+15v^{2}u_{3x}-75v_{x}v_{3x}+60uvv_{3x}\\+90v^{2}v_{3x}+5u_{xx}^{2}-20uu_{x}u_{xx}+60vv_{x}u_{xx}-45v_{xx}^{2}+90vu_{x}v_{xx}\\+90uv_{x}v_{xx}+540vv_{x}v_{xx}+30u^{2}vv_{xx}-180uv^{2}v_{xx}-90v^{3}v_{xx}\\-5u_{x}^{3}+45u_{x}v_{x}^{2}+60uvu_{x}v_{x}-180v^{2}u_{x}v_{x}+5u^{4}u_{x}\\-90u^{2}v^{2}u_{x} +45v^{4}u_{x}+180v_{x}^{3}+30u^{2}v_{x}^{2}-360uvv_{x}^{2}\\ -270v^{2}v_{x}^{2} -60u^{3}vv_{x}+180uv^{3}v_{x} \\ \\
-9v_{5x}+10vu_{4x}+25v_{x}u_{3x}+20uvu_{3x}+30v^{2}u_{3x}+15u_{x}v_{3x}+90v_{x}v_{3x}\\+15u^{2}v_{3x}+15v^{2}v_{3x}+30u_{xx}v_{xx}+50vu_{x}u_{xx}-10u^{2}vu_{xx}+50uv_{x}u_{xx}\\+60vv_{x}u_{xx}+60uv^{2}u_{xx}+30v^{3}u_{xx}+90v_{xx}^{2}+60uu_{x}v_{xx}+60vv_{x}v_{xx}\\+45u_{x}^{2}v_{x}-20uvu_{x}^{2}+60v^{2}u_{x}^{2}-10u^{2}u_{x}v_{x}+90v^{2}u_{x}v_{x}\\-20u^{3}vu_{x} +120uvu_{x}v_{x}+60uv^{3}u_{x}+15v_{x}^{3}-5u^{4}v_{x}\\+90u^{2}v^{2}v_{x} -45v^{4}v_{x} .

\end{array}
\right) .\label{new}
\end{equation}
By setting $v=0$  the well known Kupershmidt equation is an abvious reduction of system (\ref{new}):
\vspace*{-0.3cm}
\begin{equation*}
u_t=u_{5x}+5u_{x}u_{3x}+5u_{xx}^2-5u^2u_{3x}-20uu_{x}u_{xx}-5u_{x}^3+5u^4u_{x}.
\vspace*{-1cm}
\end{equation*}
 

\begin{prop}
System (\ref{new}) can be written in Hamiltonian form in not just one but two different ways: 
\begin{equation}
\left(
\begin{array}{cc} 
u_{t}\\
v_{t}
\end{array}
\right)
=
F_1 [u,v]
=
\mathrm{J}
\left(
\begin{array}{cc} 
\delta_{u}\\
\delta_{v}
\end{array}
\right)
\int \rho_{1}~ \mathrm{d}x
=\mathrm{K}
\left(
\begin{array}{cc} 
\delta_{u}\\
\delta_{v}
\end{array}
\right)
\int \rho_{-1} ~\mathrm{d}x\label{hm}
\end{equation}

with the campatible pair of Hamiltonian operators

\begin{equation*}
\mathrm{J}= \left(
\begin{array}{cc} 
 3D_x & 0  \\
 0 & D_x 
\end{array}\right),\mathrm{K} = \left(
\begin{array}{cc} 
\mathrm{K} _{1}&\mathrm{K} _{2}\\
-\mathrm{K}^{*} _{2}&\mathrm{K} _{4}
\end{array}\right) 
\end{equation*}

where

\begin{eqnarray*}
\begin{array}{ll}
K_{1}=& 2 D_x^7 +  \alpha_{1} D_x^5 + D_x^5 \alpha_{1} +  \alpha_{2} D_x^3 + D_x^3 \alpha_{2} + \alpha_{3} D_x + D_x \alpha_{3}
+4   u_{x}D_x^{-1}u_{t} +4   u_{t}D_x^{-1}u_{x}\\\\
K_{2} =&-56 D_x^6 v + D_x^5 \alpha_{4} + D_x^4 \alpha_{5} + D_x^3 \alpha_{6} + D_x^2 \alpha_{7} + D_x \alpha_{8} + \alpha_{9}
+4   u_{x}D_x^{-1}v_{t} +4   u_{t}D_x^{-1}v_{x}  \\\\
K_{4}=& -18 D_x^7 +  \alpha_{10} D_x^5 + D_x^5 \alpha_{10} +  \alpha_{11} D_x^3 + D_x^3 \alpha_{11}  + \alpha_{12} D_x + D_x \alpha_{12}
 +4   v_{x}D_x^{-1}v_{t} +4   v_{t}D_x^{-1}v_{x}  \\\\

\end{array}
\end{eqnarray*}
 
where

\begin{eqnarray*}
\begin{array}{ll}

\alpha_{1}=&6(u_{x}-u^{2}+11v^{2})\\\\

\alpha_{2}=&-16u_{3x}+20uu_{xx}-336vv_{xx}+29u_{x}^{2}-6u^{2}u_{x}-381v_{x}^{2}+9u^{4}-294u^{2}v^{2}+264uvv_{x}\\&
+45v^{4}+234v^{2}u_{x}-180v^{2}v_{x}\\\\

\alpha_{3}=&2(5u_{5x}+102vv_{4x}-25u_{x}u_{3x}+3u^{2}u_{3x}-117v^{2}u_{3x}+453v_{x}v_{3x}-72uvv_{3x}+90v^{2}v_{3x}\\&
-21u_{xx}^{2}-600vv_{x}u_{xx}+8uu_{x}u_{xx}-8u^{3}u_{xx}+264uv^{2}u_{xx}+351v_{xx}^{2}-498vu_{x}v_{xx}+540vv_{x}v_{xx}\\&-306uv_{x}v_{xx}+174u^{2}vv_{xx}-180uv^{2}v_{xx}-90v^{3}v_{xx}+6u_{x}^{3}-498u_{x}v_{x}^{2}+180v_{x}^{3}+264u^{2}v_{x}^{2}\\&
-360uvv_{x}^{2}-270v^{2}v_{x}^{2}-2u^{6}+60u^{4}v^{2}-60u^{3}vv_{x}-44u^{2}u_{x}^{2}-90u^{2}v^{4}+360u^{2}v^{2}v_{x}\\&
-6uu_{4x}+180uv^{3}v_{x}+1116uvu_{x}v_{x}+294v^{2}u_{x}^{2})\\\\

\alpha_{4}=&4(53v_{x}+28uv+42v^{2})\\\\

\alpha_{5}=&2(-165v_{xx}-192vu_{x}-212uv_{x}-312vv_{x}+32u^{2}v-168uv^{2}-72v^{3})\\\\

\alpha_{6}=&4(66v_{3x}+141vu_{xx}+165uv_{xx}+90vv_{xx}+281u_{x}v_{x}+45v_{x}^{2}-32u^{3}v-37u^{2}v_{x}-6u^{2}v^{2}\\&
+72uv^{3}-86uvu_{x}+312uvv_{x}-18v^{4}+246v^{2}u_{x}+177v^{2}v_{x})\\\\

\alpha_{7}=&2(-54v_{4x}-194vu_{3x}-264uv_{3x}-36vv_{3x}-557v_{x}u_{xx}+168uvu_{xx}-462v^{2}u_{xx}-603u_{x}v_{xx}\\&
+162v_{x}v_{xx}+57u^{2}v_{xx}-360uvv_{xx}-63v^{2}v_{xx}-288v^{3}u_{x}+184vu_{x}^{2}-1224vu_{x}v_{x}+108uv^{2}u_{x}\\&
+306uu_{x}v_{x}-180uv_{x}^{2}-246vv_{x}^{2}+148u^{3}v_{x}+24u^{3}v^{2}+336u^{2}vu_{x}+24u^{2}vv_{x}-708uv^{2}v_{x}\\&
-36v^{3}v_{x}-4u^{4}v+72uv^{4}+36v^{5})\\\\

\alpha_{8}=&2(9v_{5x}+46vu_{4x}+189v_{x}u_{3x}-80uvu_{3x}+138v^{2}u_{3x}+249u_{x}v_{3x}-90v_{x}v_{3x}-15u^{2}v_{3x}\\&
+72uvv_{3x}-15v^{2}v_{3x}+300u_{xx}v_{xx}-90v_{xx}^{2}-174uu_{x}v_{xx}+360vu_{x}v_{xx}-120vv_{x}v_{xx}-114u^{3}v_{xx}\\&
-324uv_{x}v_{xx}+126uv^{2}v_{xx}-190u_{x}^{2}v_{x}+180u_{x}v_{x}^{2}-30v_{x}^{3}+8u^{5}v+4u^{4}v_{x}+36u^{3}vu_{x}\\&
-48u^{3}vv_{x}-456u^{2}u_{x}v_{x}-192u^{2}v^{2}u_{x}-60u^{2}v^{2}v_{x}-174u^{2}vu_{xx}+108uv_{4x}-204uv_{x}u_{xx}\\&
-72uv^{5}-60uv^{3}u_{x}+72uv^{3}v_{x}-144uv^{2}u_{xx}-308uvu_{x}^{2}-168uvu_{x}v_{x}+492uvv_{x}^{2}-72v^{4}u_{x}\\&
+114v^{3}u_{xx}-84v^{2}u_{x}^{2}+648v^{2}u_{x}v_{x}+180v^{2}v_{x}^{2}-222vu_{x}u_{xx}+564vv_{x}u_{xx})\\\\

\alpha_{9}=&4(-9uv_{5x}+v_{x}u_{4x}+10uvu_{4x}+20u^{2}vu_{3x}+25uv_{x}u_{3x}+30uv^{2}u_{3x}+15u^{3}v_{3x}+15uu_{x}v_{3x}\\&
+90uv_{x}v_{3x}+15uv^{2}v_{3x}-30vv_{x}v_{3x}+5u_{x}v_{x}u_{xx}+30uu_{xx}v_{xx}+90uv_{xx}^{2}-45v_{x}^{2}v_{xx}\\&
+60u^{2}u_{x}v_{xx}+120uvv_{x}v_{xx}+90v^{2}v_{x}v_{xx}+30u^{2}vv_{x}^{2}+30vu_{x}v_{x}^{2}+30uv_{x}^{3}-4u^{5}v_{x}\\&
-20u^{4}vu_{x}-10u^{3}u_{x}v_{x}+60u^{3}v^{2}v_{x}-10u^{3}vu_{xx}+45u^{2}v_{x}u_{xx}+60u^{2}v^{3}u_{x}+60u^{2}v^{2}u_{xx}\\&
-20u^{2}vu_{x}^{2}+120u^{2}vu_{x}v_{x}+40uu_{x}^{2}v_{x}+30uv^{3}u_{xx}+60uv^{2}u_{x}^{2}+90uv^{2}u_{x}v_{x}\\&
-180uv^{2}v_{x}^{2}+50uvu_{x}u_{xx}+60uvv_{x}u_{xx}-90v^{3}v_{x}^{2}+15v^{2}v_{x}u_{xx}+180vv_{x}^{3})\\\\

\alpha_{10} =&2(9u_{x}+54v_{x}+9u^{2}+13v^{2}) \\\\

\alpha_{11} =&-36u_{3x}-288v_{3x}-72uu_{xx}+72vu_{xx}-68vv_{xx}-81u_{x}^{2}-36u_{x}v_{x}-18u^{2}u_{x}+144uvu_{x}\\&
+62v^{2}u_{x}-275v_{x}^{2}-36u^{2}v_{x}-192v^{2}v_{x}-9u^{4}+62u^{2}v^{2}-69v^{4} \\\\

\alpha_{12} =&2(9u_{5x}+90v_{5x}+18uu_{4x}-36vu_{4x}+26vv_{4x}+81u_{x}u_{3x}-54v_{x}u_{3x}+9u^{2}u_{3x}-72uvu_{3x}-51v^{2}u\\&
_{3x}+18u_{x}v_{3x}+239v_{x}v_{3x}+18u^{2}v_{3x}+96v^{2}v_{3x}+63u_{xx}^{2}+54uu_{x}u_{xx}-216vu_{x}u_{xx}-108uv_{x}u_{xx}\\&
-154vv_{x}u_{xx}+18u^{3}u_{xx}-102uv^{2}u_{xx}-60v^{3}u_{xx}+231v_{xx}^{2}-92vu_{x}v_{xx}+396vv_{x}v_{xx}-92u^{2}vv_{xx}\\&
+108v^{3}v_{xx}+18u_{x}^{3}-92v^{2}u_{x}^{2}-108u_{x}^{2}v_{x}+54u^{2}u_{x}^{2}-62u_{x}v_{x}^{2}+192v_{x}^{3}\\&
+20u^{2}v^{2}u_{x}-120uv^{3}u_{x}-308uvu_{x}v_{x}-60v^{4}u_{x}+384v^{2}v_{x}^{2}+18v^{6}+10u^{4}v^{2}\\&
-62u^{2}v_{x}^{2}-60u^{2}v^{4}) 
\end{array}
\end{eqnarray*}

\end{prop}

By straigthforward calculation it is easy to show that  the functional trivector of  linear combination $K+\lambda J$ with constant $\lambda$, vanishes independently from the value of $\lambda$ \cite{OLV}.

The first few conserved densities of the hierarchy are listed below.


\begin{eqnarray*}
\begin{array}{ll}

\rho_{-1}=&\alpha\\
\rho_{0}=&u^2 + 3v^2\\
\rho_{1}=&3uu_{4x}-81vv_{4x}-5u^{2}u_{3x}+45v^{2}u_{3x}-90uvv_{3x}-270v^{2}v_{3x}-5u^{3}u_{xx}+45uv^{2}u_{xx}+180v^{3}u_{xx}\\&
+135u^{2}vv_{xx}+45v^{3}v_{xx}+60u^{3}vv_{x}-540u^{2}v^{2}v_{x}-540uv^{3}v_{x}+u^{6}\\&-45u^{4}v^{2}+135u^{2}v^{4}-27v^{6}\\
\\

\rho_{2}=&-90uu_{6x}+7290vv_{6x}+126u^{2}u_{5x}-756v^{2}u_{5x}+13608uvv_{5x}+20412v^{2}v_{5x}+105u^{3}u_{4x}\\&-5670uv^{2}u_{4x}
-3780v^{3}u_{4x}+11340uv_{x}v_{4x}-11340u^{2}vv_{4x}-34020uv^{2}v_{4x}+16065v^{3}v_{4x}\\&-15120uvv_{x}u_{3x}-70u^{4}u_{3x}
+10080u^{2}v^{2}u_{3x}-7560uv^{3}u_{3x}-2835v^{4}u_{3x}-136080uvv_{x}v_{3x}\\&-7140u^{3}vv_{3x}+45360u^{2}v^{2}v_{3x}
+11340uv^{3}v_{3x}-17010v^{4}v_{3x}+315u^{2}u_{xx}^{2}-11340uvu_{xx}v_{xx}\\&+18900u^{2}vv_{x}u_{xx}-22680uv^{2}v_{x}u_{xx}
-252u^{5}u_{xx}+8820u^{3}v^{2}u_{xx}+13608v^{5}u_{xx}-5670u^{2}v_{xx}^{2}\\&-68040uvv_{xx}^{2}-76545v^{2}v_{xx}^{2}
-6300u^{3}v_{x}v_{xx}+136080u^{2}vv_{x}v_{xx}+34020uv^{2}v_{x}v_{xx}\\&+6930u^{4}vv_{xx}+13608v^{5}v_{xx}+7560u^{3}v^{2}v_{xx}
-22680u^{2}v^{3}v_{xx}-4410u^{4}v_{x}^{2}+15120u^{3}vv_{x}^{2}\\&-34020u^{2}v^{2}v_{x}^{2}+1512u^{5}vv_{x}
-11340u^{4}v^{2}v_{x}-68040u^{2}v^{4}v_{x}-68040uv^{5}v_{x}+30u^{8}\\&-1260u^{6}v^{2}+11340u^{2}v^{6}-2430v^{8}\\\\
\dot{.}&\\
\dot{.}&\\
\end{array}
\end{eqnarray*}

These densities suffice to write two  Magri shemes with same hamiltonian operators that one of them contains the new system proving integrability of the system (\ref{new}).

The recursion relation 

\begin{equation}
F_n
=
\mathrm{J}
\left(
\begin{array}{cc} 
\delta_{u}\\
\delta_{v}
\end{array}
\right)
\int \rho_{n}~ \mathrm{d}x
=\mathrm{K}
\left(
\begin{array}{cc} 
\delta_{u}\\
\delta_{v}
\end{array}
\right)
\int \rho_{n-2} ~\mathrm{d}x
\end{equation}

is a more general relation of (\ref{hm}) that will provide further conservation laws for n-th stage symmetry of the original system.

let us also mention that the recursion operator can be written in the form $R=KJ^{-1}$.  Starting with the basic symmetries $ \left(
\begin{array}{cc} 
u_x\\
v_x
\end{array}
\right)$ and $ \left(
\begin{array}{cc} 
u_t\\
v_t
\end{array}
\right)$ we can generate two infinite generalized symmetries  by applying

\begin{equation}
K_{n+2}=RK_n~~.
\end{equation}

\end{document}